Unusually high concentration of metals in drinking water can pose a health hazard. Twenty couples of data were taken from different locations of Kuş Cenneti Lake measuring zinc concentration in bottom water and surface water.

Table 1. Zinc Concentration (mg/l) in bottom and surface water in different regions and areas of Karacakaçan Lake

Location

Region

Area

Zinc_Bottom

Zinc_Surface

Zinc_Mean

TZ3_Mean

Loc 1

North

Area A

0.51

0.49

0.50

2.67

Loc 2

North

Area A

0.46

0.45

0.46

2.45

Loc 3

North

Area A

0.52

0.52

0.52

2.66

Loc 4

North

Area A

0.51

0.50

0.50

2.56

Loc 5

North

Area A

0.58

0.57

0.58

3.03

Loc 6

North

Area A

0.57

0.55

0.56

3.21

Loc 7

North

Area A

0.59

0.58

0.59

3.50

Loc 8

North

Area B

0.46

0.44

0.45

2.40

Loc 9

North

Area B

0.46

0.44

0.45

2.35

Loc 10

North

Area B

0.45

0.43

0.44

2.56

Loc11

East

Area B

0.41

0.40

0.41

1.33

Loc 12

East

Area B

0.41

0.39

0.40

1.56

Loc 13

East

Area B

0.41

0.39

0.40

1.23

Loc 14

East

Area B

0.47

0.45

0.46

1.12

Loc 15

East

Area C

0.41

0.39

0.40

1.23

Loc 16

East

Area C

0.37

0.36

0.37

1.10

Loc 17

East

Area C

0.41

0.39

0.40

1.20

Loc 18

East

Area C

0.40

0.39

0.39

1.10

Loc 19

East

Area C

0.39

0.37

0.38

1.12

Loc 20

East

Area C

0.46

0.44

0.45

1.30

Does the data suggest that the true average concentration in the north region is higher than that of east region? Formulating your hypothesis and providing an explanation on the reason of your choice, conduct the appropriate test and interprete your results (α=0.05)
Answer #1

From the given scenario, the null and alterate hypotheses can be stated as follows: Null hypothesis: The true average concentn 72 0.505 -0.406 0.003 0.001 + 10 10 0.099 0.020 = 4.97 P=T DIST.RT (t.df) Use Excel function T DIST RT (t.df) to get the pr= 1 + n df s m-1 m 1-1 m2 0.003 0.001 + 10 10 1 0.003 + 10-11 10 10-1 0.00000016 0.0000000118 =13 P=T DIST RT(4.97.13) =0.000